OpenStudy (anonymous):

4(6+7)=4*6+x*7 i need help... best answer will get a medal

OpenStudy (anonymous):

Hellooo :) Order of operations (or PEMDAS) says you start with what's in parentheses, then exponents, the multiplication/division, and then addition/subtraction. So first you wanna distribute the 4 on the left like: 4(6+7) is like 4 times 6 + 4 times 7 24+28=4*6+x*7 24+28=24+7x You can subtract 24 from both sides 28=7x Divide both sides by 7 x=4 Hope this helps! :)

OpenStudy (anonymous):

Thank u :-)